OpenStudy (anonymous):

-4/5 - 5/6 A. -1/30 B. -49/30 C. -9/30

OpenStudy (jamesj):

Write this as a common fraction over 5 x 6; i.e., over 30. -4/5 = -6 x 4 /30 = -24/30 -5/6 = -5 x 5 /30 = -25/30 Hence -4/5 - 5/6 = .... what?

OpenStudy (jamesj):

\[\frac{-4}{5} = \frac{-4}{5} \times \frac{6}{6} = \frac{-4 \times 6}{5 \times 6} = \frac{-24}{30}\] Likewise -5/6 = -25/30 Hence \[\frac{-4}{5} + \frac{-5}{6} = \frac{-24}{30} + \frac{-25}{30} = \frac{-24 - 25}{30} = .... what?\]

OpenStudy (jamesj):

what step are you not following?

jhonyy9 (jhonyy9):

B. is correct

OpenStudy (jamesj):

seriously jhonyy? You don't think it reasonable someone know how to add 24 and 25?